OpenStudy (anonymous):

Find the equation of the line that contains the point (-1, 2) and is perpendicular to the line 2x + 4y = -2.

OpenStudy (anonymous):

M*=2 so M=-1/2 since M=(y-y1)/(x-x1) so -1/2=(y-2)/(x+2) so 2y-4=-x-2 so 2y=2-x

OpenStudy (anonymous):

ok so what I am confused is how you figure out M

OpenStudy (anonymous):

M is the slope of the 1st line M* is the slope of the 2nd lines at which the 1st line is Perpendicular To.

OpenStudy (anonymous):

ok i understand thank you

OpenStudy (anonymous):

u wlcm